For the first two years of her mortgage Theresa was making fixed monthly payments of $1,000 per month. Her two year fixed-rate period has just ended, and now she must pay an interest rate of 6%. The amount outstanding on the mortgage is $170,000, and the mortgage will last for 28 years more. Calculate the increase in monthly repayments that she must now pay. Calculate your answer to the nearest dollar.
Answer #1

Rate = 6%

Amount outstanding (PV) = 170,000

Number of years(nper) = 28

Monthly repayment(PMT) = =PMT(6%/12,28*12,170000) PMT(rate, nper, pv, [fv], [type]) = $ 1045.71

Increase in monthly repayment = 1045.71 - 1000 = 45.71 = approx $ 46